Determine the domain of the following:
a. f(×)=3÷2×_5
b. f(×)=1÷ײ-9

Answer by Boreal(15235)

Domain of f(x)=3/(2x-5): The denominator cannot be 0, otherwise you are dividing by 0. Domain is all real numbers except x=5.
Domain of f(x)=1/(x^2-9)
Here, the denominator can't be 0, either.
x^2-9=0; x^2=9; I take the square root of x^2 to get x and the square root of 9 to get x= 3 or -3.
The domain is all reals except where x=-3 or x=3. In those instances, the denominator becomes 0.
domain comes before range in the alphabet
x comes before y in the alphabet.
domain deals with x
range deals with y